Daily Archives: June 20, 2011

Spectrometers back up

The drain valve of the air dryer went bad and needs to be replaced. A new valve has been ordered and will come soon. For time being,  the old valve is used after clean up to keep the spectrometers running. Hopefully it … Continue reading

Posted in Announcements | Comments Off

Chemistry spectrometers down

We have a problem with the lab air dryer. The spectrometers in the Chemistry NMR facility are temporarily down.

Posted in Announcements | Comments Off

I400 update

13C and 31P are ok, 1H and 19F work at reduced S/N and longer pulse widths. Probe default values have been updated but proton decoupling for 13C may be affected. 2D experiments should be run on the VXR400 and I500 … Continue reading

Posted in Announcements | Comments Off